OpenStudy (anonymous):

Find the derivative of f(x) = 5/x at x = -1

OpenStudy (xapproachesinfinity):

using the limit definition of derivative?

OpenStudy (xapproachesinfinity):

or you want it the usual way?

OpenStudy (bibby):

factor out the 5 and find the derivative of 1/x as we did before

OpenStudy (anonymous):

so it's 5(1/x), right?

OpenStudy (bibby):

correct. now we have \(\large 5\frac{d}{dx}\frac{1}{x}\)

OpenStudy (anonymous):

I honestly can't remember this next part but I'm gonna take a wild guess and go with 5(1/x^2)...?

OpenStudy (bibby):

You shouldn't be guessing. Why is that the case? What is the derivative of \(\frac{1}{x}\)?

OpenStudy (xapproachesinfinity):

oops no guessing hehe

OpenStudy (anonymous):

x^-1?

OpenStudy (bibby):

\(\large \frac{1}{x}=x^{-1}\)

OpenStudy (anonymous):

okay, so we plug it into the power rule now?

OpenStudy (bibby):

correct.

OpenStudy (anonymous):

f(x)=x^-1 -> -x^(-1-1) -> -x^-2

OpenStudy (bibby):

nice. yeah. so now bring it back to the original equation. \(\large 5\frac{d}{dx}\frac{1}{x}=5*\frac{1}{x^2}=\frac{5}{x^2}\)

OpenStudy (anonymous):

and then plug in -1 for x so it's 5/1 or just 5?

OpenStudy (bibby):

looks right

OpenStudy (anonymous):

It's finally making sense! Thanks :D

OpenStudy (bibby):

yay, I'm so proud ;~;